# A First-Order Gradient Approach for the Connectivity Analysis of Markov Chains

### v2

@@ Abstract
Weighted graphs are commonly used to model various complex systems, including social networks, power grids, transportation networks, and biological systems. In many applications, the connectivity of these networks can be expressed through the Mean First Passage Times (MFPTs) of a Markov chain modeling a random walker on the graph. In this paper, we generalize the network metrics based on Markov chains’ MFPTs and extend them to networks affected by uncertainty, in which edges may fail and hence not be present according to a pre-determined stochastic model. To find optimally connected Markov chains, we present a parameterization-free method for optimizing the MFPTs of the Markov chain. More specifically, we present an efficient Simultaneous Perturbation Stochastic Approximation (SPSA) algorithm in the context of Markov chain optimization. The proposed algorithm is suitable for both fixed and random networks. Using various numerical experiments, we demonstrate scalability compared to established benchmarks. Importantly, our algorithm finds an optimal solution without requiring prior knowledge of edge failure probabilities, allowing for an online optimization approach.
@@ I Introduction
Weighted graphs are widely used to describe complex systems, such as social networks, power grids, transportation networks, and biological systems. In many applications, it is often useful to study the dynamical properties of these networked systems by analyzing the behavior of random walks on them. Such a random walk can be specified as a Markov chain by explicitly giving all transition probabilities between each pair of nodes. Mean first passage times (MFPTs) characterize the connectivity properties of the given network, as they describe the expected time for the random walker starting at a given node to reach another specific node for the first time, for a review see [1, 2]. MFPTs are studied in various fields where phenomena can be modeled with Markovian dynamics, for instance, in biology [3] or chemistry [4, 5].
@@
In this paper, we focus on analyzing and minimizing MFPTs of Markov chains. Several metrics for network connectivity and robustness have been proposed in the literature that can be written as weighted combinations of mean first passage times, among which the well-known effective graph resistance [6] or Kirchhoff index [7], the DW-Kirchhoff index [8], and the Kemeny constant [9].
@@
A large body of literature studies how to optimally design or modify edge weights and, hence, the network structure to improve such metrics; see, e.g., [10, 11]. In view of the above interpretation, this class of problems can be reformulated as finding the Markov chain that minimizes a specific function of the MFPTs. Generally, the resulting problem is non-convex, but restricting the set of solutions to reversible Markov chains can turn the problem into a convex one. For example, the minimization of the Kemeny constant can be made convex by constraining the stationary distribution of the Markov chain and consequently can be solved using semi-definite programming [11, 12]. However, the restriction to reversibility has a severe impact on the achievable objective value (see [12] for an example), which motivates our research into a gradient-based optimization method that does not rely on convexity and thus allows one to drop the reversibility condition on the network.
@@
In the present paper, we deviate from the standard approaches in the literature by formulating and solving more general and possibly non-convex network design problems. In particular, we allow for more general MFPT-based objective functions. Moreover, we extend the existing literature by considering settings in which network edges can fail with some (possibly correlated) probability. For instance, this can be of particular interest when modeling power grids where transmission lines may fail due to climatic disasters (wildfires, floods, storms, etc.). Therefore, in the rest of the paper, we distinguish between the fixed support case, in which the set of edges of the network is fixed, and the random support case, where there is uncertainty in the support of the graph, i.e., edges fail stochastically.
@@
  • If the considered graph is Hamiltonian, we derive explicit bounds for the sum of the MFPTs if we assume reversibility in Proposition III.2, and we rigorously show that we can obtain a strictly better solution by relaxing the reversibility assumption found in Proposition III.1.
  • We modify the Simultaneous Perturbation Stochastic Approximation (SPSA) algorithm [13] in the context of Markov chain optimization to solve possibly non-convex network optimization problems, both in the fixed support setting (see Theorem IV.3), as well as in the random support setting (see Theorem V.2). Concurrently, we allow for constraining the stationary distribution of the Markov chain. We show that our method outperforms a benchmark in terms of scalability and can seamlessly integrate an online optimization approach. Our relatively simple yet powerful extension of SPSA underscores the versatility of first-order methods in studying network connectivity.
@@
The paper is organized as follows. Section II introduces the notation and some preliminaries. In Section III, we describe the network connectivity optimization problem with the generalized MFPT metric as objective in the fixed support setting and provide conditions for its convexity. In the same section, for some special graphs of interest, we also derive explicit bounds for the so-called “price of reversibility”. We derive closed-form expressions for the derivative of MFPTs with respect to the parameters of the Markov chain in Section IV, and leverage them to introduce an SPSA-type method for the optimization of Markov chains. In Section V, we consider the problem in the random support setting, explicitly list sufficient conditions for convexity, and explain how to modify our algorithm accordingly. In Section VI, we demonstrate SPSA’s scalability and its handling of correlated edge presence in the random support setting. We present an application of SPSA in a surveillance context in Section VII. Using our SPSA, we obtain a solution that, leveraging directionality and irreversibility, strictly improves reversible solutions. We conclude with a discussion of further research in Section VIII.
@@ II Notation and preliminaries
We consider a finite, directed, and simple graph $G=(V,E)$, with set of nodes $V=\{1,\ldots,N\}$ and set of directed edges $E\subset V\times V$. We denote by $E_{i}$ the set of edges leaving node $i$, i.e., $E_{i}=\{\ell\in E:\ell=(i,j),j\in V\}$. Given $G=(V,E)$, we consider the vector $x\in[0,\infty)^{|E|}$ whose elements are the edge weights of $G$, that is, $x_{\ell}=x_{(i,j)}$ is the weight of edge $\ell=(i,j)\in E$. We denote the vector of weights of the edges leaving node $i\in V$ by
@@
Throughout the paper, we specifically consider weights that induce a Markov chain on a graph $G$. More formally, let $\mathcal{P}=\mathcal{P}(G)$ be the set of stochastic matrices associated with $G$, i.e., the collection of $|V|\times|V|$ matrices $P$ such that $P_{ij}\geq 0$ if $(i,j)\in E$ and $P_{ij}=0$ otherwise. Then, for $x\in[0,\infty)^{|E|}$ and $\sum_{\ell\in E_{i}}x_{\ell}>0$, for all $i\in V$, we define the Markov chain associated with $x$ through
$P(x)=P(x,G)=\sum_{\ell=(i,j)\in E}\frac{x_{\ell}}{\sum_{\ell\in E_{i}}x_{\ell}}e_{i}e_{j}^{\top}\in\mathcal{P},$ (3)
where $e_{i}\in\mathbb{R}^{|V|}$ is the vector with a 1 in the $i$th position and zeros elsewhere. For our analysis, we work with the following constraint sets:
@@
$P(x)=P(x,G)=\sum_{\ell=(i,j)\in E}x_{\ell}e_{i}e_{j}^{\top}\in\mathcal{P}.$ (4)
Throughout the paper, we will refer to a general stochastic matrix associated with $G$ as $P$ if there is no explicit dependence on $x$.
The Markov chain associated with a given stochastic matrix $P\in\mathcal{P}(G)$ describes a random walk on the graph $G$. More formally, if $X_{n}\in V$ denotes the node of the network at which the random walker resides at time $n\in\mathbb{N}$, we assume that for every $i,j\in V$, the transition probabilities from node $i$ to node $j$ only depend on the current node and are given by the corresponding element of the matrix $P$, i.e.,
@@
Throughout the paper, we assume that $G$ is strongly connected. This assumption implies that any $P(x)$ is irreducible whenever $x\in\mathcal{X}_{\varepsilon}$. In an irreducible Markov chain, all nodes in the graph “communicate”, i.e., for all $(i,j)\in V\times V$, there exists an index $k\in\mathbb{N}$, such that $P_{ij}^{k}>0$. An irreducible $P(x)$ has a unique stationary distribution $\pi=\pi(P(x))$, for which $\pi\geq 0$, $\sum\pi=1$, and $\pi P(x)=\pi$ (where following the standard convention, we treat $\pi$ as a row vector). We denote
$\mathcal{X}^{\hat{\pi}}=\{x\in\mathcal{X}:\pi(P(x))=\hat{\pi}\},\text{ and }\mathcal{X}^{\pi}_{\varepsilon}=\mathcal{X}^{\hat{\pi}}\cap\mathcal{X}_{\varepsilon},$
as sets inducing Markov chains $\mathcal{P}^{\pi}=\{P\in\mathcal{P}:\pi(P)=\hat{\pi}\}$ with fixed stationary distribution $\hat{\pi}$ via (4).
The value $\pi_{i}$ of the stationary distribution of $P$ can be interpreted as the long-run fraction of visits to node $i$ by the random walker. In particular, it a.s. holds that
$\pi_{ij}$ $=$ $\lim_{n\to\infty}\frac{1}{n}\sum_{m=1}^{n}\text{Prob}(X_{m}=j~{}|~{}X_{0}=i)$
$=$ $\lim_{n\to\infty}\frac{1}{n}\sum_{m=1}^{n}(P^{m})_{ij}=\lim_{n\to\infty}\frac{1}{n}\sum_{m=1}^{n}{\bf 1}(X_{m}=i),$
for $X_{0}=j$ and ${\bf 1}(\cdot)$ being the indicator function. The matrix
$\Pi=(\pi_{ij})_{i,j\in V}$
is called the ergodic projector of $P$. If $P$ is irreducible (but not necessarily aperiodic), the limits $\pi_{ij}$ are independent of the initial node $j$ (i.e., $\pi_{i}=\pi_{ij}$ for all $j\in V$) so that $\Pi$ has equal rows, all equal to the unique stationary distribution $\pi=(\pi_{i}:i\in V)$ of $P$.
For a Markov chain $P$, let $M\in\mathbb{N}^{N\times N}$ be the matrix of the mean first passage times, where for every $i,j\in V$
@@
Given $\Pi$ and $D$, the mean first passage time matrix $M$ can be obtained using the following closed-form expression [9]:
$M=(I-D+{\bar{1}}{\bar{1}}^{\top}\cdot\textup{dg}(D))\cdot\textup{dg}(\Pi)^{-1},$ (6)
@@ A Markov chain $P$ is reversible if it satisfies the following condition
A frequently used quantity in the design of undirected networks is the effective graph resistance or Kirchhoff index. For a graph $(G,x)$, $x\in\mathcal{X}^{\mathrm{sym}}$, the effective graph resistance $R_{tot}(x)$ is defined using the effective resistance matrix $R(x)=R(G,x)\in\mathbb{R}_{\geq 0}^{N\times N}$ as follows
$R_{tot}(x)=R_{tot}(G,x)=\sum_{i<j}R_{ij}(x),$ (8)
@@ III Markov chain connectivity on fixed graphs
In this paper, our goal is to find optimally connected Markov chains by minimizing $S(P,C)$. From now on, we will treat $C$ as a scalar matrix; however, it is important to note that our optimization techniques also apply to cases where $C$ varies with $P$, for example, when optimizing the graph’s Kemeny constant.
@@ III-A Optimizing Markov chain connectivity
Let us formulate the problem of designing an optimally connected Markov chain $P\in\mathcal{P}$ on a graph $G$ as
@@
$\min_{x\in{\cal X}^{(\mathrm{sym},1)}}$ $S(P(x),C)$ (15)
Then, the reversibility condition (R) is satisfied for $P(x)$ via (3) for all strongly connected $(G,x)$ and the problem in (15) is convex. Moreover, if $C$ is a matrix with identical entries $C_{ij}=c>0$, then (15) is a strictly convex problem.
@@
As argued earlier, we can model all $P$ that satisfy condition (R) using $\mathcal{X}^{(\mathrm{sym},1)}$, which is clearly convex in view of its linear constraints. To show the convexity of $S(P(x),C)$, we use the convexity of all pairwise resistance distances $R_{ij}(x)$ in any undirected weighted graph $(G,x)$ shown in [10]. It follows that for any symmetric matrix $C\in\mathbb{R}_{\geq 0}^{N\times N}$ we can rewrite
@@
We do not need to explicitly model the constraint that $(G,x)$ is strongly connected in Theorem III.1, since any graph, where $C_{ij}>0$ and $i$ and $j$ are in separate components, has $S(P(x),C)=\infty$ (due to $M_{ij}=\infty$).
@@ III-B The price of reversibility
In the following, we consider a specific case that allows for comparing analytical reversible and non-reversible solutions to (12). To that end, consider a directed graph $G$ that is Hamiltonian, which means that $G$ admits a path, called Hamiltonian cycle, that starts and ends at the same node, visiting all other nodes in the graph exactly once. The following theorem states that if $G$ is a Hamiltonian graph and if we assume unit weights $C=C^{\bar{1}}$ for the mean first passage times in (10), then the optimal solution of (12) yields a Hamiltonian cycle on $G$.
@@
Let $G=(V,E)$ be a directed Hamiltonian graph. Then, the Markov chain
@@
corresponds to a Hamiltonian cycle on $G$, that is, $\hat{P}$ is irreducible, and each row has one entry equal to $1$ and all other entries equal to 00. Furthermore, its optimal value equals
@@
See Appendix -D. ∎
If the graph $G$ admits a Hamiltonian cycle and we assume reversibility, the following lower and upper bounds hold for the optimal Markov chain $\hat{P}$.
@@
Let $G=(V,E)$ be a directed Hamiltonian graph with $N\geq 3$ nodes. Consider the optimization problem (18) with $P$ satisfying (R) as an additional constraint;
$\check{P}=\operatorname*{arg\,min}_{P\in\mathcal{P}^{\mathrm{rev}}}\ S(P,C^{\bar{1}})$ (20)

### v4

# A First-Order Gradient Approach for the Connectivity Optimization of Markov Chains

@@ Abstract
Graphs are commonly used to model various complex systems, including social networks, power grids, transportation networks, and biological systems. In many applications, the connectivity of these networks can be expressed through the Mean First Passage Times (MFPTs) of a Markov chain modeling a random walker on the graph. In this paper, we generalize the network metrics based on Markov chains’ MFPTs and extend them to networks affected by uncertainty, in which edges may fail and hence not be present according to a pre-determined stochastic model. To find optimally connected Markov chains, we present a parameterization-free method for optimizing the MFPTs of the Markov chain. More specifically, we present an efficient Simultaneous Perturbation Stochastic Approximation (SPSA) algorithm in the context of Markov chain optimization. The proposed algorithm is suitable for both fixed and random networks. Using various numerical experiments, we demonstrate scalability compared to established benchmarks. Importantly, our algorithm finds an optimal solution without requiring prior knowledge of edge failure probabilities, allowing for an online optimization approach.
@@ I Introduction
Graphs are widely used to describe complex systems, such as social networks, power grids, transportation networks, and biological systems. In many applications, it is often useful to study the dynamical properties of these networked systems by analyzing the behavior of random walks on them. Such a random walk can be specified as a Markov chain by explicitly giving all transition probabilities between each pair of nodes. Mean first passage times (MFPTs) characterize the connectivity properties of the given network, as they describe the expected time for the random walker starting at a given node to reach another specific node for the first time, for a review, see [1, 2]. MFPTs are studied in various fields where phenomena can be modeled with Markovian dynamics, for instance, in biology [3] or chemistry [4, 5].
@@
In this paper, we focus on analyzing and minimizing mean first passage times (MFPTs) in Markov chains. MFPTs serve as building blocks for several important metrics of network connectivity and robustness. Examples include the effective graph resistance or Kirchhoff index [6, 7], the DW-Kirchhoff index [8], and the Kemeny constant [9], all of which can be expressed as weighted combinations of MFPTs.
A central theme in the literature is how to optimally design or modify edge weights to improve such metrics; see, e.g., [10, 11]. In this context, optimizing edge weights can be equivalently interpreted as designing an optimal exploration policy over the graph. Indeed, since edge weights define the transition probabilities of the associated random walk, they implicitly determine how the graph is traversed. From this perspective, the problem becomes one of policy optimization, where the goal is to identify the Markov chain that minimizes a specified MFPT-based performance criterion.
@@
The resulting problem is generally non-convex, but restricting the set of solutions to reversible Markov chains can turn the problem into a convex one. For example, the minimization of the Kemeny constant can be made convex by constraining the stationary distribution of the Markov chain and consequently can be solved using semi-definite programming [11, 12]. However, the restriction to reversibility has a severe impact on the achievable objective value (see [12] for an example), which motivates our research into a gradient-based optimization method that does not rely on convexity and thus allows one to drop the reversibility condition on the network.
@@
In the present paper, we deviate from the standard approaches in the literature by formulating and solving more general and possibly non-convex optimization problems for network connectivity. In particular, we allow for more general MFPT-based objective functions. Moreover, we extend the existing literature by considering settings in which network edges can fail with some (possibly correlated) probability. For instance, this can be of particular interest when modeling power grids where transmission lines may fail due to climatic disasters (wildfires, floods, storms, etc.). Therefore, in the rest of the paper, we distinguish between the fixed support case, in which the set of edges of the network is fixed, and the random support case, where there is uncertainty in the support of the graph, i.e., edges fail stochastically.
@@
  • If the considered graph is Hamiltonian, we derive explicit bounds for the sum of the MFPTs if we assume reversibility in Proposition III.2, and we show that we can obtain a strictly better solution by relaxing the reversibility assumption.
  • We modify the Simultaneous Perturbation Stochastic Approximation (SPSA) algorithm [13, 14] specifically for the context of Markov chain optimization to solve possibly non-convex policy optimization problems, both in the fixed support setting (see Theorem IV.3), as well as in the random support setting (see Theorem V.2). A key feature of our approach is a tailored perturbation and projection scheme that guarantees all iterates remain valid Markov chains, thereby respecting the hard constraints intrinsic to the problem of Markov chain optimization. Concurrently, we allow for constraining the stationary distribution of the Markov chain. We show that our method outperforms a benchmark in terms of scalability and can seamlessly integrate an online optimization approach. Our relatively simple yet powerful extension of SPSA underscores the versatility of first-order methods in studying network connectivity.
@@
The paper is organized as follows. Section II introduces the notation and some preliminaries. In Section III, we describe the Markov chain optimization problem with the generalized MFPT metric as objective in the fixed support setting and provide conditions for its convexity. In the same section, for some special graphs of interest, we also derive explicit bounds for the so-called “price of reversibility”. We derive closed-form expressions for the derivative of MFPTs with respect to the parameters of the Markov chain in Section IV, and leverage them to introduce an SPSA-type method for the optimization of Markov chains. In Section V, we consider the problem in the random support setting, explicitly list sufficient conditions for convexity, and explain how to modify our algorithm accordingly. In Section VI, we demonstrate SPSA’s scalability and its handling of correlated edge presence in the random support setting. We present an application of SPSA in a surveillance context in Section VII. Using our SPSA, we obtain a solution that, leveraging directionality and irreversibility, strictly improves reversible solutions. We conclude with a discussion of further research in Section VIII.
@@ II Notation and preliminaries
We consider a finite, directed, and simple graph $\mathscr{G}=(V,E)$, with set of nodes $V=\{1,\ldots,N\}$ and set of directed edges $E\subset V\times V$. We denote by $E_{i}$ the set of edges leaving node $i$, i.e., $E_{i}=\{\ell\in E:\ell=(i,j),j\in V\}$. Given $\mathscr{G}=(V,E)$, we consider the vector $x\in[0,\infty)^{|E|}$ whose elements are the edge weights of $\mathscr{G}$, that is, $x_{\ell}=x_{(i,j)}$ is the weight of edge $\ell=(i,j)\in E$. We denote the vector of weights of the edges leaving node $i\in V$ by
@@
Throughout the paper, we specifically consider weights that induce a Markov chain on a graph $\mathscr{G}$. More formally, let $\mathcal{P}=\mathcal{P}(\mathscr{G})$ be the set of stochastic matrices associated with $\mathscr{G}$, i.e., the collection of $N\times N$ matrices $P$ such that $P_{ij}\geq 0$ if $(i,j)\in E$ and $P_{ij}=0$ otherwise. Then, for $x\in[0,\infty)^{|E|}$ and $\sum_{\ell\in E_{i}}x_{\ell}>0$, for all $i\in V$, we define the Markov chain associated with $x$ through
$P(x)=P(x,\mathscr{G})=\sum_{\ell=(i,j)\in E}\frac{x_{\ell}}{\sum_{\ell\in E_{i}}x_{\ell}}e_{i}e_{j}^{\top}\in\mathcal{P},$ (3)
where $e_{i}\in\mathbb{R}^{N}$ is the vector with a 1 in the $i$th position and zeros elsewhere. For our analysis, we work with the following constraint sets:
@@
$P(x)=P(x,\mathscr{G})=\sum_{\ell=(i,j)\in E}x_{\ell}e_{i}e_{j}^{\top}\in\mathcal{P}.$ (4)
Throughout the paper, we will refer to a general stochastic matrix associated with $\mathscr{G}$ as $P$ if there is no explicit dependence on $x$.
The Markov chain associated with a given stochastic matrix $P\in\mathcal{P}(\mathscr{G})$ describes a random walk on the graph $\mathscr{G}$. More formally, if $X_{n}\in V$ denotes the node of the network at which the random walker resides at time $n\in\mathbb{N}$, we assume that for every $i,j\in V$, the transition probabilities from node $i$ to node $j$ only depend on the current node and are given by the corresponding element of the matrix $P$, i.e.,
@@
Throughout the paper, we assume that $\mathscr{G}$ is strongly connected. This assumption implies that any $P(x)$ is irreducible whenever $x\in\mathcal{X}_{\varepsilon}$. In an irreducible Markov chain, all nodes in the graph “communicate”, i.e., for all $(i,j)\in V\times V$, there exists an index $k\in\mathbb{N}$, such that $P_{ij}^{k}>0$. An irreducible and finite $P(x)$ is positive recurrent and so has a unique stationary distribution $\pi=\pi(P(x))$, such that
$\pi>0,\ \sum_{i=1}^{N}\pi_{i}=1,\text{ and }\pi P(x)=\pi.$
Throughout the paper, we adopt the convention that $\pi(P(x))$ (or simply $\pi$ when $P$ is clear from context) is a row vector and always refers to this unique stationary distribution. Moreover, we denote
[…]
as sets inducing Markov chains $\mathcal{P}^{\hat{\pi}}=\{P\in\mathcal{P}:\pi(P)=\hat{\pi}\}$ with fixed stationary distribution $\hat{\pi}$ via (4).
The value $\pi_{j}$ of the stationary distribution of $P$ can be interpreted as the long-run fraction of visits to node $j$ by the random walker. A convenient way to formalize this is via the ergodic projector of $P$, that is, the matrix $\Pi:=(\Pi_{ij})_{i,j\in V},$ whose entries are defined as
$\Pi_{ij}:=\lim_{n\to\infty}\frac{1}{n}\sum_{m=1}^{n}\text{Prob}(X_{m}=j~|~X_{0}=i).$
In particular, if $P$ is irreducible (but not necessarily aperiodic), for any $i\in V$ it holds a.s. that
$\Pi_{ij}=\lim_{n\to\infty}\frac{1}{n}\sum_{m=1}^{n}{\bf 1}(X_{m}=j),$
where ${\bf 1}(\cdot)$ denotes the indicator function. In other words, the limits $\Pi_{ij}$ are independent of the initial node $i$ (i.e., $\pi_{j}=\Pi_{ij}$, for all $j\in V$), so that $\Pi$ has equal rows, all equal to the unique stationary distribution $\pi=(\pi_{j}:j\in V)$ of $P$.
For a Markov chain $P$, let $M\in\mathbb{R}_{\geq 0}^{N\times N}$ be the matrix of the mean first passage times, where for every $i,j\in V$
@@
Given $\Pi$ and $D$, the mean first passage time matrix $M$ can be obtained using the following closed-form expression [9, 15]:
$M=(I-D+{\bar{1}}{\bar{1}}^{\top}\textup{dg}(D))\cdot\textup{dg}(\Pi)^{-1},$ (6)
@@ A Markov chain $P$ is reversible if it satisfies the following condition
A frequently used quantity to study network connectivity is the effective graph resistance or Kirchhoff index. For a graph $(\mathscr{G},x)$, $x\in\mathcal{X}^{\mathrm{sym}}$, the effective graph resistance $R_{tot}(x)$ is defined using the effective resistance matrix $R(x)=R(\mathscr{G},x)\in\mathbb{R}_{\geq 0}^{N\times N}$ as follows
$R_{tot}(x)=R_{tot}(\mathscr{G},x)=\sum_{i<j}R_{ij}(x),$ (7)
@@ III Markov chain connectivity on fixed graphs
In this paper, our goal is to find optimally connected Markov chains by minimizing $S(P,C)$. From now on, we will treat $C$ as a scalar matrix; however, it is important to note that our optimization techniques also apply to cases where $C$ varies with $P$ (as long as $C(P)$ is three times differentiable), for example, when optimizing the graph’s Kemeny constant.
@@ III-A Optimizing Markov chain connectivity
Let us formulate the problem of designing an optimally connected Markov chain $P\in\mathcal{P}$ on a graph $\mathscr{G}$ as
@@
$\min_{x\in{\cal X}^{(\mathrm{sym},1)}}$ $S(P(x),C).$ (14)
Then, the reversibility condition (R) is satisfied for $P(x)$ via (3) for all strongly connected $(\mathscr{G},x)$ and the problem in (14) is convex. Moreover, if $C$ is a matrix with identical entries $C_{ij}=c>0$, then (14) is a strictly convex problem.
@@
As argued earlier, we can model all $P$ that satisfy condition (R) using $\mathcal{X}^{(\mathrm{sym},1)}$, which is clearly convex in view of its linear constraints. To show the convexity of $S(P(x),C)$, we use the convexity of all pairwise resistance distances $R_{ij}(x)$ in any undirected weighted graph $(\mathscr{G},x)$ shown in [10]. It follows that for any symmetric matrix $C\in\mathbb{R}_{\geq 0}^{N\times N}$ we can rewrite
@@
We do not need to explicitly model the constraint that $(\mathscr{G},x)$ is strongly connected in Theorem III.1, since any graph, where $C_{ij}>0$ and $i$ and $j$ are in separate components, has $S(P(x),C)=\infty$ (due to $M_{ij}=\infty$).
@@ III-B The price of reversibility
In the following, we consider a specific case that allows for comparing analytical reversible and non-reversible solutions to (11). To that end, consider a directed graph $\mathscr{G}$ that is Hamiltonian, which means that $\mathscr{G}$ admits a path, called Hamiltonian cycle, that starts and ends at the same node, visiting all other nodes in the graph exactly once. The following theorem states that if $\mathscr{G}$ is a Hamiltonian graph and if we assume unit weights $C=C^{\bar{1}}$ for the mean first passage times in (9), then the optimal solution of (11) yields a Hamiltonian cycle on $\mathscr{G}$.
@@
Let $\mathscr{G}=(V,E)$ be a directed Hamiltonian graph. Then, the Markov chain
@@
corresponds to a Hamiltonian cycle on $\mathscr{G}$, that is, $\hat{P}$ is irreducible, and each row has one entry equal to $1$ and all other entries equal to 0. Furthermore, its optimal value equals
@@
The proof straightforwardly follows from [24, Theorem 2], see Appendix -D. ∎
If the graph $\mathscr{G}$ admits a Hamiltonian cycle and we assume reversibility, the following lower and upper bounds hold for the optimal Markov chain $\hat{P}$.
@@
Let $\mathscr{G}=(V,E)$ be a directed Hamiltonian graph with $N\geq 3$ nodes. Consider the optimization problem (17) with $P$ satisfying (R) as an additional constraint;
$\check{P}=\operatorname*{arg\,min}_{P\in\mathcal{P}^{\mathrm{rev}}}\ S(P,C^{\bar{1}}).$ (19)